# Toward thermoelectric characterization of (nano)materials by in situ transmission electron microscopy

### v1

@@ 2.4 Electrical properties
In addition to the thermoelectric characterization, the chips also facilitate an in situ TEM measurement of the electrical properties of the materials. The electrical conductivity can be obtained from the measured resistance and the sample geometry. The data obtained from the devices at room temperature (=0) are summarized in SI12.
As the measurements were conducted in a two-probe setup, the contribution of contact resistance needs to be considered, which was approximately 135 for the employed chips. For specimens with dimensions typical for TEM investigations, the contact resistance is only a negligible part of the total device resistance. For example, the resistance of the MLC NT with a diameter of 100 nm and a length of 2.8 m is 10 k and in this case, the contact resistance is only a small fraction of the total resistance. The derived conductivity amounts to 3.7 104 Sm-1, a value which agrees with previous measurements [22, 24]. Similarly, the two thinned CCO MLC devices show resistances of 104 k (CCO ) and 80 k (CCO ), which lead to conductivities of 7 Sm-1 and 5 Sm-1, respectively. These values are of the same order of magnitude, but slightly smaller when comparing to bulk measurements [7]. This could be attributed to the fact that pores, absent in the studied microscopic devices but frequent in the bulk (see SI7), significantly contribute to the resistivity of the CCO material.
For the Si (890 ), the unthinned CCO (430 ) and especially for the Mo device (186 ), which are thick compared to standard TEM specimen thicknesses, the contact resistance is a large contribution to the measured total resistance. In case of the Mo device, a reliable measurement of the conductivity is thus not possible, which would require thinning of the sample to thicknesses below 100 nm over an extended area to reach larger resistance values. For the unthinned CCO device, the high conductivity ( = 5.8 103 Sm-1) is explained by the contribution from the Pt cover. Finally, for the Si device, the calculated resistivity of 1.3 10-3 m is considerably lower than the indicated resistivity of the employed Si wafer (0.1 m). This is caused by Ga implantation during FIB milling (approximately 0.5 at%, EDX analysis in SI3), which acts as additional p-doping of the Si, and possibly as well by remaining Pt on the device surface.
@@ 2.5 Thermoelectric properties
All the studied devices clearly show the induction of a thermovoltage, which increases in magnitude with the temperature gradient created along the material by the differential heating device. The experiments thus prove the possibility to measure thermoelectric properties of materials with the proposed setup. However, for a quantitative analysis, the temperature gradient needs to be known in dependence of the applied heating current . We performed COMSOL simulations of the temperature distribution under vacuum within the silicon nitride membrane for different and specimen parameters. Figure 4c and f depict two temperature maps simulated for = 5 mA and assuming a small thermal conductivity of the device material (1 Wm-1K-1). The maps illustrate the effect of one important device parameter: the length of the hole milled in the membrane between the contact pads. A short hole, similar in size to the devices analyzed so far, barely affects the temperature distribution in the silicon nitride membrane (left part in Figure 4e), leading to a small temperature difference between the pads. Only a long hole across the entire membrane forms an effective heat barrier and yields a higher temperature (gradient) for the same , which then merely depends on the thermal conductivity and the geometry of the specimen.
To test this simulation result, we milled a longer hole in the membrane of the CCO (66 m) and the Mo device (77 m) and repeated the measurement of the voltage offset in dependence of . SI13 shows the SEM images of the devices with long hole. The result is plotted in Figure 4d in comparison with the same devices with a short hole and the CCO device, for which a long hole (65 m) was milled directly upon preparation. The measured voltage offset strongly increases by almost two orders of magnitude for the CCO device, while the sign changes for the Mo device, which only shows a small change in magnitude. This striking difference is explained by two aspects: the contribution of the Pt contacts and the large difference in thermal conductivity between the two devices. With regard to the first aspect, the contribution of the Pt contacts has been ignored so far. Pt has an absolute Seebeck coefficient (-5 VK-1) with the same magnitude as Mo (5 VK-1), but with inverted sign. In the device with a short hole, the temperature gradient along both contact paths is largely similar and they cancel each other out almost entirely and the contribution from the Mo device is dominant. In contrast, the temperature gradient along the Pt contacts is asymmetric and their contribution outweighs the contribution of the Mo device for a long hole. Regarding the second aspect, the thermal conductivity is much higher for the metallic Mo (139 Wm-1K-1) than for the CCO (3 Wm-1K-1) [7]. Therefore, the temperature gradient along the CCO device strongly increases, leading to the observed strong increase in the voltage offset.
@@
The COMSOL simulations indicate that only a hole milled across the whole membrane provides a full prevention of heat flux along the membrane of the chip. However, when opening the hole over the whole membrane, the mechanical tension created along the investigated material that links both sides of the membrane is too large, causing the fracture of the device (SI14).
@@
To get a rough estimation of the value of the Seebeck coefficient of the CCO devices, a value for the temperature gradient needs to be obtained. Therefore, the heater temperature can be calculated from the measured resistance and the calibrated linear temperature coefficient of Pt (SI15), which results in a temperature of approximately 125 and 140°C for the highest applied heating currents for the CCO devices (Figure 4e). Although the COMSOL simulations don’t show a quantitative agreement with this estimated temperate, they were used to estimate the ratio between heater temperature and the actual temperature gradient across the device to 3:1 (SI16). Under these assumptions, the Seebeck coefficients of the CCO and the CCO devices are calculated to 135 and 125 VK-1. Although the values correspond well with bulk values [7], they have to be seen as a very rough estimation and the real values could differ considerably.
@@ 2.6 Toward fully quantitative thermoelectric characterization
Despite the strong increase in the induced thermovoltage and the good agreement of the estimated Seebeck coefficients, a reliable quantification still requires an exact knowledge of the temperature gradient between the contact pads, which cannot be determined from the employed chip without assuming a value for the thermal conductivity and calculations/simulations of the heat transfer of the device. Due to this limitation, the used chip design is not suitable to fully exploit the advantage of performing thermoelectric measurements by in situ TEM, which would allow studying the interplay between material properties and structure/composition on the local scale.
To reach the possibility of a full quantification, Figure 5 shows a sketch of a device design, which could facilitate such measurements. The design would require an in situ TEM holder with at least eight electrical contacts. While two contacts would be used for electrical characterization of the device (S-, S+), the remaining six would be designated to two fully symmetrical heating elements. For both, two are used as current supply (H1,2-, H1,2+) and the third one as temperature sensor (H1,2 sense). Temperature measurement is performed in a three-point setup requiring identical dimensions of the two current paths for the heating elements.
@@
The design shown in Figure 5 would have to be placed on top of a silicon nitride membrane. With such a design, it would be possible to determine the temperature gradient and thus to quantify the Seebeck coefficient as well as the conductivity of the studied devices. For example, for the CCO material it would be possible to disentangle the contributions from crystal orientation, grain boundaries or even pores. It is noted that for the employed electron microscope, TEM holders are only available with four contacts. The complete development of a new holder with eight contacts is out of the scope of this work.
@@ 3 Conclusion
To conclude, our experiments prove that a thermoelectric characterization of micro- and nanosized devices by in situ transmission electron microscopy (TEM) is possible. A thermovoltage is clearly induced in all studied devices and the sign corresponds to the sign of the materials’ Seebeck coefficient. The results indicate that the temperature gradient is linearly proportional to the deposited heat energy and that an effective heat barrier should be generated between hot and cold side of the device to obtain a defined temperature gradient along the device. We propose an improved chip design and measurement setup for in situ TEM holders with eight electrical contacts, which should provide the combination of full quantitative thermoelectric characterization with TEM and electron spectroscopies. Such studies will allow to investigate in detail, e.g., the effect of grain boundaries and dopants, in principle down to the single-atom level, on the thermoelectric properties of bulk and nanomaterials. The in situ studies can be further enriched by additional, complementary microscopy techniques performed on the identical device, such as Raman microscopy [24].
Experimental Section
MEMS-based chip fabrication
Thermoelectric in situ chips were fabricated by MEMS technology starting with a four-inch 100 Si wafer (380 m thickness) coated on both sides with 1 m of low-stress SiNx and following this recipe:
1.
Spin coating of Ti35 ESX neg resist (Microchemicals) on the back side with 3000 rounds per minute (rpm).
2.
3 min bake at 100 °C on hot plate.
3.
Optical lithography defining windows for later membrane creation with 175 mJcm-2 and hard contact.
4.
3 min bake at 130 °C on hot plate.
5.
Flood exposure with 520 mJcm-2.
6.
Development for 95s in a 1:1 mixture of H2O : AZ Developer (Microchemicals).
7.
Reactive ion etching (20 sccm SF6, 0.19 mbar, 200 W, 15 min) of the back side to remove the SiNx.
8.
Spin coating of AZ5214E (Microchemicals) with 6000 rpm on the top side.
9.
Waiting time of 10 min.
10.
50s bake at 110 °C on hot plate.
11.
Optical lithography of front side defining heating element and conductive paths with 26 mJcm-2 and soft contact.
12.
1 min bake at 120 °C on hot plate.
13.
Flood exposure with 390 mJcm-2.
14.
Development for 35s in AZ 726 MiF (Merck).
15.
Sputtering of 5 nm Ti + 150 nm Pt (AJA Orion 5 UHV).
16.
Lift-off in acetone assisted by ultrasound.
17.
Membrane creation by wet etching in KOH at 80 °C with a holder for front-side protection of the Pt contacts.
18.
Cutting of wafer in individual chips by a wafer saw.
Electron microscopy
The whole chip is designed to fit in a DENSsolutions Wildfire holder with 4 pins for a ThermoFisherScientific/FEI instrument, which allows TEM analysis and electrical characterization of the chip at the same time. TEM, SAED and scanning (S)TEM analysis have been conducted in two aberration-corrected microscopes (Thermo Fisher Scientific Titan3 (image-corrected) and Titan low-base (probe-corrected)) operated at 300 keV electron energy. Energy-dispersive X-ray spectroscopy (EDX) was performed in STEM mode with an Ultim X-MaxN 100TLE detector (Oxford Instruments).
Bulk CCO material preparation
The CCO bulk material was prepared as described in [7] by attrition milling (2h at 600 rpm) starting from SrCO3, CaCO3 and CoO. The mixture was calcinated at 850 °C during 1 h, cold uniaxially pressed into 25 mm diameter disks under 250 MPa, and subjected to hot-uniaxial pressing at 900 °C and 51 MPa for 1 h.
In situ device preparation
A dual-beam instrument (Thermo Fisher Scientific Helios 650) with SEM and focused ion beam (FIB) has been utilized to prepare the in situ TEM devices. For bulk materials, FIB-based preparation methods for in situ studies have been published [23] and their first step is the preparation of a conventional TEM lamella: a vertical slice of material with 1 m thickness, 12 m width and a depth of approximately 5 m is taken out from the bulk material and transferred to a TEM copper grid. In a second step, the slice is flipped and transferred to the contact pads of the in situ chip, where a hole has been milled previously between the pads by FIB. The lamella can be thinned prior to transfer to the chip and/or directly on the chip by using a sample holder with inclined surface. Here, we did a combination of both. At first we thinned the sheet on the Cu grid down to 1 m thickness over the entire width and polished the surface that will be in contact to the chip to guarantee a good contact. Except for the CCO device, which was transferred without addtional pre-thinning, we further thinned the central part of the sheet from one side down to 700 nm forming a kind of bridge. This shape facilitates the thinning after the transfer of the flipped sheet to the chip as it moves the incidence of the focused ion beam away from the surface. The final thinning of the central part and the removal of the Pt cover from the TEM lamella preparation was done using an inclined sample holder that allows milling with the FIB parallel to the chips’ surface. SI6 and SI8 describe this preparation for the two CCO samples.
For the MLC NT sample, the preparation needs to be different and we employed a recently developed support-based method [24]. SI10 shows the preparation of the MLC NT device, which consists in first preparing a conventional TEM sample by drop-casting the nanomaterial on a holey silicon nitride chip and the selection of a suitable NT by TEM. Using FIB, the NT and the support are transferred to the in situ chip, where finally the silicon nitride is milled away to leave the NT as unique bridge between the contacts.
Electrical characterization
Supporting Information
Supporting Information is available from the author.

### v2

@@ 3.3 Electrical properties
In addition to the thermoelectric characterization, the chips also facilitate an in-situ TEM measurement of the electrical properties of the materials. The electrical conductivity can be obtained from the measured resistance and the sample geometry. The data obtained from the devices at room temperature (=0) are summarized in SI9.
As the measurements were conducted in a two-probe setup, the contribution of contact resistance needs to be considered, which was approximately 135 for the employed chips. For semiconducting specimens with dimensions typical for TEM investigations, the contact resistance is only a small part of the total device resistance. For example, the two thinned CCO MLC devices show resistances of 10 k (CCO ) and 80 k (CCO ), which lead to conductivities of 71 Sm-1 and 54 Sm-1, respectively. These values are one order of magnitude higher when comparing to bulk measurements [7]. This could be attributed to the fact that pores, absent in the studied microscopic devices but frequent in the bulk (see SI7), significantly contribute to the resistivity of the CCO material.
For the Si (890 ), the unthinned CCO (430 ) and especially for the metallic Mo device (186 ), which are rather thick compared to standard TEM specimen thicknesses, the contact resistance is a large contribution to the measured total resistance. In case of the Mo device, a reliable measurement of the conductivity is thus not possible, which would require a strong decrease of the sample size. For the unthinned CCO device, the high conductivity ( = 5.8 103 Sm-1) is explained by the contribution from the Pt cover. Finally, for the Si device, the calculated resistivity of 1.3 10-3 m is considerably lower than the indicated resistivity of the employed Si wafer (0.1 m). This is caused by Ga implantation during FIB milling (approximately 0.5 at%, EDX analysis in SI5), which acts as additional p-doping of the Si, and possibly as well by remaining Pt on the device surface.
Finally, the resistance of the MLC NT, which possesses a diameter of 100 nm and a length of 2.8 m, is 10 k and the contact resistance is only a small fraction of the total resistance. The derived conductivity amounts to 3.7 104 Sm-1, a value which agrees with previous measurements [23, 22]. This confirms that using the recently developed transfer process for nanomaterials [22] allows to study them in a pristine state.
@@ 3.4 Thermoelectric properties
All the studied devices clearly show the induction of a thermovoltage, which increases in magnitude with the temperature gradient created along the material by the differential heating device. The experiments thus prove the possibility to measure thermoelectric properties of materials with the proposed setup. However, for a quantitative analysis, the temperature gradient needs to be known in dependence of the applied heating current . We performed COMSOL simulations of the temperature distribution under vacuum within the silicon nitride membrane for different and specimen parameters. Figure 4c and f depict two temperature maps simulated for = 5 mA and assuming a small thermal conductivity (1 Wm-1K-1) of the specimen, on the order of magnitude of the CCO material. The maps illustrate the effect of one important device parameter: the length of the hole milled in the membrane between the contact pads. A short hole, similar in size to the devices analyzed so far, barely affects the temperature distribution in the silicon nitride membrane (left part in Figure 4e), leading to a small temperature difference between the pads. Only a long hole across the entire membrane forms an effective heat barrier and yields a higher temperature (gradient) for the same , which then merely depends on the thermal conductivity and the geometry of the specimen.
To test this simulation result, we milled a longer hole in the membrane of the CCO (66 m) and the Mo device (77 m) and repeated the measurement of the voltage offset in dependence of . SI10 shows the SEM images of the devices with long hole. The result is plotted in Figure 4d in comparison with the same devices with a short hole and the CCO device, for which a long hole (65 m) was milled directly upon preparation. The measured voltage offset strongly increases by almost two orders of magnitude for the CCO device, while the sign changes for the Mo device, which only shows a small change in magnitude. This striking difference is explained by two aspects: the contribution of the Pt contacts and the large difference in thermal conductivity between the two devices. With regard to the first aspect, the contribution of the Pt contacts has been ignored so far. Pt has an absolute Seebeck coefficient (-5 VK-1) with the same magnitude as Mo (5 VK-1), but with inverted sign. In the device with a short hole, the temperature gradient along both contact paths is largely similar and they cancel each other out almost entirely and the contribution from the Mo device is dominant. In contrast, the temperature gradient along the Pt contacts is asymmetric and their contribution outweighs the contribution of the Mo device for a long hole. Regarding the second aspect, the thermal conductivity is much higher for the metallic Mo (139 Wm-1K-1) than for the CCO (3 Wm-1K-1) [7]. Therefore, the temperature gradient along the CCO device strongly increases, leading to the observed strong increase in the voltage offset.
@@
The COMSOL simulations indicate that only a hole milled across the whole membrane provides a full prevention of heat flux along the membrane of the chip. However, when opening the hole over the whole membrane, the mechanical tension created along the investigated material that links both sides of the membrane is too large, causing the fracture of the device (SI11).
@@
To get a rough quantitative estimation of the Seebeck coefficient of the CCO specimens, a value for the temperature gradient needs to be obtained. Therefore, the heater temperature can be calculated from the measured resistance and the calibrated linear temperature coefficient of Pt (SI12), which results in a temperature of approximately 125 and 140°C for the highest applied heating currents for the CCO devices (Figure 4e). Although the COMSOL simulations don’t show a quantitative agreement with this estimated temperate, they were used to estimate the ratio between heater temperature and the actual temperature gradient across the device to 3:1 (SI13). Under these assumptions, the Seebeck coefficients of the CCO and the CCO devices are calculated to 135 and 125 VK-1. Although the values correspond well with bulk values [7], they have to be seen as a very rough estimation and the real values could differ considerably. The difference between CCO and devices could indicate a variation of the Seebeck coefficient with the lattice direction for the CCO material and a possible influence of grain boundaries. Further studies with a fully quantitative setup (see section 3.6) are required to provide more insights in this relationship and to determine the role of specific types of grain boundaries.
3.5 Dynamic thermoelectric characterization during in-situ crystallization of an amorphous Ge thin film
The experiments and analysis presented so far were performed in a static way without inducing structural changes in the specimen during the in-situ TEM studies. Such a static analysis can provide important information on thermoelectricity as it allows to correlate the thermoelectric properties with the structure and composition of the entire studied specimen at high resolution down to the atomic scale. For example, by preparing devices with specific microstructures or defects, it would be possible to determine the contribution of a specific grain boundary on the Seebeck coefficient. However, an in-situ approach typically refers to the ability to track a dynamic evolution of a specimen within the microscope. Such a dynamic evolution could be generated by increasing the temperature using the heating element, but the temperatures reachable with the used chip design are not high enough to induce significant changes in the analyzed specimens. Another approach is to employ an electrical current to induce a structural modification in the specimen mainly by Joule heating [21].
As an example for such a dynamic in-situ thermoelectric characterization, we tracked the induced thermovoltage during the crystallization process of an amorphous Ge (aGe) thin film upon application of elevated current densities. Therefore, a first in-situ device with long hole (85 m) was prepared from an aGe thin film with a thickness of 36 nm (SI14). Its thermoelectric characterization is depicted by black diamond symbols in Figure 5a and reveals a high positive induced voltage offset of almost 10 mV at = 5 mA, even higher than the one measured for the CCO materials (Figure 4d). However, Ge rapidly forms a passivating oxidation layer after being exposed to air that generates an electrically insulating barrier (SI14) and impedes a controlled flow of the high electrical currents required for inducing an in-situ crystallization of the thin film.
In order to prevent this passivation, an aGe thin film was sandwiched between two amorphous C (aC) layers in a single evaporation step without venting the chamber. Figure 5b and c show a SEM and HAADF STEM image of the prepared device and a SAED pattern (inset in Figure 5c) confirms the amorphous nature of both aC and aGe layers. A radial profile of the SAED pattern is shown in Figure 5d by a blue line revealing two broad peaks centered around 3 and 5.4 nm-1. A low-loss EEL spectrum (blue line in Figure 5e) confirms the presence of both Ge, reflected by the bulk plasmon at 16 eV and the Ge-M edge,[31] and aC, which is visible from the bulk plasmon at 25 eV [21]. A thermoelectric characterization of the device, which was prepared directly with a long hole (66 m), is depicted by blue crosses in Figure 5a and reveals a positive induced voltage with a magnitude ten times smaller compared to the pristine, passivated aGe thin film.
To reduce a possible impact of the aC layers on the measurement, an O2/Ar plasma cleaning step of 1 min was performed to remove most of the carbon from the specimen. The effect of the cleaning is reflected in a slight decrease of scattered intensity in the SAED pattern (red line in Figure 5d) and by an almost complete removal of the aC bulk plasmon in the low-loss EEL spectrum (red line in Figure 5e) when compared to the initial state (blue lines). A measurement of the voltage offset showed a slight increase of the induced voltage offset (red circles in Figure 5a), despite the increase in device resistance from 42 k to 64 k due to the removal of a part of the aC layer. As spectroscopic techniques indicated the presence of remaining C in the specimen (SI15), a second plasma cleaning step of 30s was performed. Indeed, the low-loss EEL spectrum reveals a further decrease in intensity in the region of the aC bulk plasmon (yellow line in Figure 5e), although spectroscopic analysis showed still a considerable presence of C (SI15). A large portion of the aGe thin film crystallized during this step, probably induced by an accidental electrical discharge. This crystallization into nano-grains is seen from the BFTEM image and the inset SAED pattern in Figure 5g. The border between crystallized and remaining amorphous areas is indicated by yellow lines. This crystallization translated to a more than three-fold increase of the induced voltage offset (yellow pluses in Figure 5a).
Following this unintended initial crystallization, we applied a series of double current sweeps with increasing maximum current up to 240 A to continue the crystallization process in a controlled way. The sweeps are shown in Figure 5f, where solid and dashed lines depict the up- and down-sweeps, respectively, and the color changes from blue to red with increasing maximum current. The calculated resistance at the maximum current of each sweep is plotted by black crosses and is observed to be constant (73 k) up to an applied current of 13 A before starting to decrease due to the induced Joule heating and a related increase of conductivity of the semiconducting Ge. Above an applied current of approximately 70 A, the up- and down-sweeps start to differ considerably, indicating a structural modification of the specimen induced by the electrical current. This modification is linked to, firstly, a growth of the total size of the crystallized area and, secondly, a growth of the grain size in the central area. The first trend is observed from the corresponding BFTEM image (Figure 5h), which reveals that almost the entire suspended thin film has crystallized and only small amorphous regions remain, as indicated by purple lines. The second aspect is visible from the decreasing number of reflection spots in the SAED pattern (inset in Figure 5h), which correspond to fcc Ge (Crystallography Open Database #7101739) as marked in the purple radial profile in Figure 5d. The crystal structure is as well confirmed by HAADF STEM imaging (SI15). The dynamic increment in crystallized area can be observed from the series of BFTEM images in supplementary movie S1 taken after each of the double current sweeps. This controlled increase in crystallization leads to a decrease in (static) resistance down to 18 k, which goes in hand with a further rise of the induced voltage offset (purple squares in Figure 5a).
The crystallization of the specimen is clearly seen by the comparison of Raman spectra acquired from the device after the experiment and from as-prepared thin films suspended on standard Cu TEM grids (Figure 5i). While the aGe specimens (blue and black line in Figure 5i) show broad peaks at 85 and 265 cm-1 corresponding to the longitudinal (LO) and transversal optical (TO) modes, a sharp TO mode, shifted to 295 cm-1, is the dominant contribution in the spectrum of the crystallized in-situ specimen (purple line), which also exhibits the 2TO mode. These features correspond well to Ge reference spectra from literature [32]. A contribution of the aC is seen by a very broad peak in the range between 1000 and 1700 cm-1 for the sandwiched aGe specimen and in strongly reduced intensity for the in-situ specimen.
This in-situ experiment demonstrates the possibilities of performing dynamic thermoelectric characterizations combined with TEM characterization, which allows the correlation of the microstructure, crystallinity and composition of a specimen at high spatial resolution with its thermoelectric properties. The experiment also can serve to illustrate that these measurements can be conducted with minimum and negligible contamination if the recently developed transfer method is employed [22]. In EDX spectra acquired from the aGe specimens, no Pt contamination could be found and Ga implantation was limited to the very edge of the specimen (SI15). The EDX analysis indicates a percentage of less than 0.5 at% of Ga and less than 0.1 at% of Pt in the central part of the thin film. A better estimation of a maximum amount of implantation is given by the measured electrical resistances assuming Ga as p-doping element in the Ge thin film. Neglecting the contribution from the remaining aC layer and the edges, the final resistance of 18 k leads to a resistivity of 0.34 mm, which gives an upper estimate of Ga concentration of 30 ppm, following the relation in [33]. The concentration is expected to be even lower, due to the contribution from the aC and the fact that nano-sized objects typically exhibit higher conductivities compared to bulk values.
A comparison of the measured voltage offsets allows the following conclusions. Generally, the positive value for agrees with the positive sign of the Seebeck coefficient of Ge [34]. The magnitude of the measured for the crystallized aGe device is comparable to the CCO devices (Seebeck coefficient of approximately 150 VK-1, Figure 4d), which agrees with the expected magnitude for Ge from literature [34]. The fact that the removal of a big part of the aC layer from the sandwiched aGe specimen only leads to a minimum increase of (blue crosses and red circles in Figure 5a) suggests that the Seebeck coefficient is dominated by the aGe even if it is encapsulated by aC, which has a significantly higher electrical conductivity but a negligible Seebeck coefficient. This could be attributed to the amorphous nature of the entire specimen, which could cause a strong intermixing of charge carriers between the layers of the different materials. The small increase in could be (partially) caused by an increased temperature gradient along the specimen due to a reduced thermal conductivity after removal of the aC layer.
is observed to increase with the degree of crystallinity, which is attributed to the increase in electrical conductivity of the Ge layer, allowing for a higher mobility of the charge carriers due to a reduced number of sources of electron scattering within the thin film. This finding as well agrees with literature [34]. Further in-situ TEM studies could shed light on the dependence of the Seebeck coefficient on specific crystal orientations and/or the number and type of grain boundaries. Finally, the highest is observed for the passivated aGe thin film, which shows that the insulating passivation layer has a strong impact on the Seebeck coefficient of the device. A possible explanation could be that the insulating GeOx surface layer on its own contributes to the measured thermovoltage. Another contribution could be from the reduced thermal conductivity of the passivated aGe compared to the sandwiched specimen, which would cause a difference in temperature gradient.
@@ 3.6 Toward fully quantitative thermoelectric characterization
Despite the strong induced thermovoltage, its dynamic increment with increasing crystallization of an aGe thin film and the good agreement of the estimated Seebeck coefficients for the CCO materials, a reliable quantification still requires an exact knowledge of the temperature gradient between the contact pads, which cannot be determined from the employed chip without assuming a value for the thermal conductivity and calculations/simulations of the heat transfer of the device. Due to this limitation, the used chip design is not suitable to fully exploit the advantage of performing thermoelectric measurements by in-situ TEM, which would allow studying the interplay between material properties and structure/composition on the local scale.
To reach the possibility of a full quantification, Figure 6 shows a sketch of a device design, which could facilitate such measurements. The design would require an in-situ TEM holder with at least eight electrical contacts. While two contacts would be used for two-probe electrical characterization of the device (S-, S+), the remaining six would be designated to two fully symmetrical heating elements. For both, two are used as current supply (H1,2-, H1,2+) and the third one as temperature sensor (H1,2 sense). Temperature measurement is performed in a three-point setup requiring identical dimensions of the two current paths for the heating elements.
@@
The design shown in Figure 6 would have to be placed on top of a silicon nitride membrane. With such a design, it would be possible to determine the temperature gradient and thus to quantify the Seebeck coefficient as well as the conductivity of the studied specimens. For example, for the CCO material it would be possible to disentangle the contributions from crystal orientation, grain boundaries or even pores. The design would also allow to perform a controlled in-situ heating by setting both heating elements to the same temperature and thus follow, e.g., the thermoelectric properties of a material during annealing.
@@ 4 Conclusion
To conclude, our experiments prove that a thermoelectric characterization of micro- and nanosized devices by in-situ transmission electron microscopy (TEM) is possible. A thermovoltage is clearly induced in all studied devices and the sign corresponds to the sign of the materials’ Seebeck coefficient. The results indicate that the temperature gradient is linearly proportional to the deposited heat energy and that an effective heat barrier should be generated between hot and cold side of the device to obtain a defined temperature gradient along the device. We propose an improved chip design and measurement setup for in-situ TEM holders with eight electrical contacts, which should provide the combination of full quantitative thermoelectric characterization, including thermal and electrical conductivities, with TEM and electron spectroscopies. The detailed analysis of the devices also shows that, especially for nanomaterials, specimens can be prepared with negligible artifacts allowing their investigation in a pristine state.
Such studies will allow to investigate in detail, e.g., the effect of crystal directions in asymmetric materials, grain boundaries and dopants, in principle down to the single-atom level, on the thermoelectric properties of bulk and nanomaterials. The experiments can be conducted in a static or dynamic way. Static characterization is yielded by preparing various devices with specific structural properties such as exhibiting a certain type of grain boundary or a determined crystal direction with respect to the heat and current flow as exemplary shown for two CCO devices. Dynamical studies allow the tracking of thermoelectrical properties during, e.g., in-situ annealing of thermoelectric materials, as exemplary demonstrated by the in-situ crystallization of an amorphous Ge thin film by Joule heating. The in-situ studies can be further enriched by additional, complementary microscopy techniques performed on the identical device, such as Raman microscopy.
The in-situ TEM technique lends itself on the one hand for fundamental studies as the entire device can be analyzed in depth and thus be well compared to first-principles calculations that take into account the defects present in the specimen. On the other hand, the evolution of the properties of a thermoelectric material during annealing allows a more applied approach of the technique.
Supplementary material
Supplementary material is available under the following link: https://doi.org/10.1016/j.ultramic.2024.114071